# DistriFusion: Distributed Parallel Inference for High-Resolution Diffusion Models

### 29 Feb 2024 v1

@@ 1 Introduction
Recent efforts to accelerate diffusion model inference have mainly focused on two approaches: reducing sampling steps [58, 62, 37, 21, 71, 33, 34, 74] and optimizing neural network inference [24, 26, 27]. As computational resources grow rapidly, leveraging multiple GPUs to speed up inference is appealing. For example, in natural language processing (NLP), large language models have successfully harnessed tensor parallelism across GPUs, significantly reducing latency. However, for diffusion models, multiple GPUs are usually only used for batch inference. When generating a single image, typically only one GPU is involved (Figure 2(a)). Techniques like tensor parallelism are less suitable for diffusion models, as communication costs outweigh savings from distributed computation. Thus, even when multiple GPUs are available, they cannot be effectively exploited to further accelerate single-image generation. This motivates the development of a method that can utilize multiple GPUs to speed up single-image generation with diffusion models.
@@ 2 Related Work
Parallelism. Existing work has explored various parallelism strategies to accelerate the training and inference of large language models (LLMs), including data, pipeline [15, 39, 28], tensor [17, 40, 72, 73, 79], and zero-redundancy parallelism [48, 51, 52, 78]. Tensor parallelism in particular has been widely adopted for accelerating LLMs [29], as LLM inference tends to be memory-bound. In such scenarios, the communication overhead introduced by tensor parallelism is relatively minor compared to the substantial latency benefits brought by increased memory bandwidth. However, the situation differs for diffusion models, which are compute-bound. For diffusion models, the communication overhead from tensor parallelism becomes a significant factor, overshadowing the actual computation time. As a result, only data parallelism has been used thus far for diffusion model serving, which provides no latency improvements. The only exception is ParaDiGMS [60], which uses Picard iteration to run multiple steps in parallel. However, this sampler tends to waste much computation, and the generated results exhibit significant deviation from the original diffusion model. Our method is based on patch parallelism, which distributes the computation across multiple devices by splitting the input into small patches. Compared to tensor parallelism, such a scheme has superior independence and reduced communication demands. Additionally, it favors the use of AllGather over AllReduce for data interaction, significantly lowering overhead (see Section 5.3 for the full comparisons). Drawing inspiration from the success of asynchronous communication in parallel computing [68], we further reuse the features from the previous step as context for current step to overlap communication and computation, called displaced patch parallelism. This represents the first parallelism strategy tailored to the sequential characteristics of diffusion models while avoiding the heavy communication costs of traditional techniques like tensor parallelism.
@@ 5.3 Ablation Study
Communication cost. In Table 2, we benchmark our latency with synchronous tensor parallelism (Sync. TP) and synchronous patch parallelism (Sync. PP), and report the corresponding communication amounts. Compared to TP, PP eliminates the need for communication within cross-attention and linear layers. For convolutional layers, communication is only required at the patch boundaries, which represent a minimal portion of the entire tensor. Moreover, PP utilizes AllGather over AllReduce, leading to lower communication demands and no additional use of computing resources. Therefore, PP requires $60\%$ fewer communication amounts and is 1.62.1×1.6$\sim$ 2.1$\times$1.6 ∼ 2.1 × faster than TP. We also include a theoretical PP baseline without any communication (No Comm.) to demonstrate the communication overhead in Sync. PP and DistriFusion. Compared to Sync. PP, DistriFusion further cuts such overhead by over $50\%$. The remaining overhead mainly comes from our current usage of NVIDIA Collective Communication Library (NCCL) for asynchronous communication. NCCL kernels use SMs (the computing resources on GPUs), which will slow down the overlapped computation. Using remote memory access can bypass this issue and close the performance gap.

### 07 Mar 2024 v2

@@ 1 Introduction
Recent efforts to accelerate diffusion model inference have mainly focused on two approaches: reducing sampling steps [58, 62, 37, 21, 71, 33, 34, 74] and optimizing neural network inference [24, 26, 27]. As computational resources grow rapidly, leveraging multiple GPUs to speed up inference is appealing. For example, in natural language processing (NLP), large language models have successfully harnessed tensor parallelism across GPUs, significantly reducing latency. However, for diffusion models, multiple GPUs are usually only used for batch inference. When generating a single image, typically only one GPU is involved (Figure 2(a)). Techniques like tensor parallelism are less suitable for diffusion models due to the large activation size, as communication costs outweigh savings from distributed computation. Thus, even when multiple GPUs are available, they cannot be effectively exploited to further accelerate single-image generation. This motivates the development of a method that can utilize multiple GPUs to speed up single-image generation with diffusion models.
@@ 2 Related Work
Parallelism. Existing work has explored various parallelism strategies to accelerate the training and inference of large language models (LLMs), including data, pipeline [15, 39, 28], tensor [17, 40, 72, 73, 79], and zero-redundancy parallelism [48, 51, 52, 78]. Tensor parallelism, in particular, has been widely adopted for accelerating LLMs [29], which are characterized by their substantial model sizes, whereas their activation sizes are relatively small. In such scenarios, the communication overhead introduced by tensor parallelism is relatively minor compared to the substantial latency benefits brought by increased memory bandwidth. However, the situation differs for diffusion models, which are generally smaller than LLMs but are often bottlenecked by the large activation size due to the spatial dimensions, especially when generating high-resolution content. The communication overhead from tensor parallelism becomes a significant factor, overshadowing the actual computation time. As a result, only data parallelism has been used thus far for diffusion model serving, which provides no latency improvements. The only exception is ParaDiGMS [60], which uses Picard iteration to run multiple steps in parallel. However, this sampler tends to waste much computation, and the generated results exhibit significant deviation from the original diffusion model. Our method is based on patch parallelism, which distributes the computation across multiple devices by splitting the input into small patches. Compared to tensor parallelism, such a scheme has superior independence and reduced communication demands. Additionally, it favors the use of AllGather over AllReduce for data interaction, significantly lowering overhead (see Section 5.3 for the full comparisons). Drawing inspiration from the success of asynchronous communication in parallel computing [68], we further reuse the features from the previous step as context for current step to overlap communication and computation, called displaced patch parallelism. This represents the first parallelism strategy tailored to the sequential characteristics of diffusion models while avoiding the heavy communication costs of traditional techniques like tensor parallelism.
@@ 5.3 Ablation Study
Compare to tensor parallelism. In Table 2, we benchmark our latency with synchronous tensor parallelism (Sync. TP) and synchronous patch parallelism (Sync. PP), and report the corresponding communication amounts. Compared to TP, PP has better independence, which eliminates the need for communication within cross-attention and linear layers. For convolutional layers, communication is only required at the patch boundaries, which represent a minimal portion of the entire tensor. Moreover, PP utilizes AllGather over AllReduce, leading to lower communication demands and no additional use of computing resources. Therefore, PP requires $60\%$ fewer communication amounts and is 1.62.1×1.6$\sim$ 2.1$\times$1.6 ∼ 2.1 × faster than TP, making it a more efficient approach for deploying diffusion models. We also include a theoretical PP baseline without any communication (No Comm.) to demonstrate the communication overhead in Sync. PP and DistriFusion. Compared to Sync. PP, DistriFusion further cuts such overhead by over $50\%$. The remaining overhead mainly comes from our current usage of NVIDIA Collective Communication Library (NCCL) for asynchronous communication. NCCL kernels use SMs (the computing resources on GPUs), which will slow down the overlapped computation. Using remote memory access can bypass this issue and close the performance gap.